The length of a rectangle is five times its width.
If the perimeter of the rectangle is 72m, find its area.

Answer by JBarnum(2146)

given/known
l=5w
p=2(l+w)
a=lw
.
72=2((5w)+w)
36=6w
6=w
.
a=(5w)w=5w^2
a=5(6)^2
a=5(36)
a=180
.